# Aleph-Alpha-GermanWeb: Improving German-language LLM pre-training with model-based data curation and synthetic data generation

### v2

@@ 2.3 Quality classification
Following prior work on quality classification for English-language web data [46], we categorised documents into five quality buckets – high, medium-high, medium, medium-low, and low – by leveraging the strengths of several imperfect quality classifiers to produce a more accurate outcome. Specifically, we created several individual quality classifiers based on fastText [7] and Bidirectional Encoder Representations from Transformers (BERT) [11] models. In the following, we describe (1) our two grammar classifiers, (2) our two educational quality classifiers, (3) our two instruction classifiers and (4) the way in which we ensembled all classifiers and grouped the data into the five quality buckets according to the classifiers’ scores.
@@ 2.3.2 Educational quality
Inspired by successful demonstrations for English with FineWeb [38] and Nemotron-CC [46] (Apachee 2.0), we created educational quality classifiers to curate high-quality German web data by training a classifier on scores given by an LLM-as-a-judge. The intuition behind this is to identify web documents that are more likely to be educational, informative, and useful for model training, as opposed to spam, low-quality, or otherwise unhelpful content.
@@
For each document, we calculated a combined educational quality score by taking the minimum over the three criteria rated by the LLM-as-a-judge. This is because we observed that the judge would often miss poor quality in one criterium, but rarely in all three. We then used these educational quality scores as the training signal for fastText and BERT quality classification models. We trained one BERT model tasked to predict the educational quality scores given the first 512 tokens of the document’s text. We also trained a binary classifier based on fastText.
@@ 2.3.3 Instruction
Building on Messmer et al. [31], we trained binary classifiers using both fastText and BERT to distinguish between chat/instruction-style data (high quality) and other types of documents (low quality). To create a balanced training dataset, we created two subsets of 80,000 documents each. The low-quality subset consisted of randomly selected documents from the German FineWeb2 corpus, while the high-quality subset comprised randomly sampled instances from the Aya collection, Include Base-44, OpenAssistant2, and MMLU. To create samples from MMLU, we first concatenated a question and the respective answers. For OpenAssistant-2, we concatenated all samples that belong to the same conversation. We used 95% of the data for training and the remaining 5% for validation.
@@ 2.3.4 Ensembling
To split the data into different quality buckets, we use the rules in Table 3 to first allocate overall quality points to each document.
Model Condition Points Initialise all documents 0 BERT (edu. quality) Predicted score = 5 +3 fastText (edu. quality) Predicted “high quality” with >99% confidence +2 BERT (grammar) Predicted “high quality” +3 fastText (grammar) Predicted “high quality” with >99% confidence +2 BERT (instruction) High quality probability is in top 15% of the distribution +6 fastText (instruction) High quality probability is in top 15% of the distribution +4
Using the overall quality scores, we categorised documents into five quality buckets based on the overall quality scores shown in Table 4.
Quality Bucket Score Range Quality Distribution of FineWeb2
@@
Medium-high $9-11$ 15.0% 9.9%
Medium $5-8$ 36.3% 28.7%
Medium-low $3-4$ 15.5% 21.7%
Low $<3$ 21.0% 31.6%
Table 4: Document quality classifications decided by overall quality score.
@@ 3 Experiments
We trained and evaluated a one billion parameter (1B) Llama-style transformer model and an eight billion parameter (8B) hierarchical autoregressive transformer (HAT) model on our dataset to assess its quality compared to FineWeb2.
@@ 3.1 Evaluations
The Multilingual Massive Multitask Language Understanding dataset [18], is a benchmark designed to evaluate the performance of LLMs across multiple languages and disciplines. It extends the original MMLU benchmark by translating its test set into 14 languages—including German—using professional human translators to ensure accuracy. The dataset encompasses 57 subjects ranging from elementary-level topics to advanced professional fields such as law, physics, history, and computer science. The German portion contains 14,000 samples.
The following three popular LLM benchmarks were translated from English to German [40].
@@
The AI2 Reasoning Challenge (ARC) [9] is a benchmark dataset developed by the Allen Institute for AI to evaluate the reasoning capabilities of artificial intelligence models. It comprises 7,787 multiple-choice science questions sourced from standardized tests designed for students in grades three through nine. The dataset is divided into two subsets: ARC-Easy and ARC-Challenge; here we evaluate the translation of ARC-Easy.
@@
The HellaSwag benchmark dataset [54] is designed to evaluate the commonsense reasoning abilities of AI models, particularly in the context of sentence completion tasks. It comprises approximately 70,000 multiple-choice questions from diverse sources, including instructional videos and articles from platforms like WikiHow and ActivityNet. Each question presents a context followed by four possible sentence completions, one of which is correct.
@@
The TruthfulQA dataset [26] is a benchmark designed to evaluate the truthfulness of language models when generating answers to questions. It comprises 817 questions across 38 categories, including health, law, finance, and politics. The questions are crafted to challenge models with scenarios where humans might hold incorrect beliefs or misconceptions, aiming to assess whether models can avoid generating false answers learned from imitating human texts.
Figure 1: Average accuracies (MMMLU, ARC-Easy, HellaSwag; single- & five-shot) of 1B Llama-style models trained on $\sim$ 84 billion tokens from different datasets. All subsets of Aleph-Alpha-GermanWeb – marked with an asterisk (*) – outperform FineWeb2. For comparison, we also include EPFML High. Here, 1,000 training steps equates to approximately 2.1 billion tokens. Individual benchmark results are provided in App.  D.1.
@@
We trained and evaluated our filtered Common Crawl (Filtered CC), synthetic, and our FineWeb2 High bucket, and compared these to random sampling German FineWeb2 and the top 10% of German FineWeb2 documents classified by Messmer et al. [31] (EPFML High). The results are shown in Figure 1. We chose to exclude TruthfulQA from the 1B parameter model evaluations, as we found that it did not give a clear signal and suspect that this size of the model may be incapable of the task (cf. §3.3 for results of an 8B model).
@@
To evaluate our dataset on a different architecture and a larger scale, we pre-trained 8B hierarchical autoregressive transformer (HAT) models [34] (see Appendix C.2 for training details). With this we leverage recent advances demonstrating that jointly training the tokenizer and model can be a promising alternative [37], aligning with the philosophy of end-to-end learning  [48].
An additional benefit of using an end-to-end, tokenizer-free architecture like HAT is that it debiases comparisons between datasets—comparisons which are a central focus of this paper. Tokenizers have a significant impact on the efficacy of model training [2, 52, 49]. Typically they are trained on large text corpora, which makes models using those tokenizers more effective on datasets that resemble the tokenizer’s training distribution. As a result, using a tokenizer-based model can introduce bias when comparing the effectiveness of different pre-training datasets, favouring those models that align more closely with the tokenizer’s original training data [30]. To avoid this bias, we use a HAT model, which does not rely on any pre-trained tokenizer and instead performs fixed, dataset-agnostic, byte-level string splitting.
@@
When pre-training LLMs, web data is often mixed from different languages and with data from alternative, curated sources [16], e.g., English and German data might be mixed with additional non-web data from books and encyclopedias. To test the performance of our single highest-performing subset, we compared our synthetic data to mix of 50% FineWeb2 and 50% high-quality curated datasets [^4] In particular: the German National Library, European Union Parliament transcripts, European Patent Office applications, German Text Archive, Project Gutenberg books, JRC-Acquis, NewsAPI, Open Legal Data, Open Web Math, peS2o, PhilPapers, Wikibooks, Wikinews, Wikpedia, Wikisource, and Wikivoyage.. Figure 2 (left) shows our synthetic data outperformed this mixture on our tested benchmarks.
@@
Figure 2: Accuracies of 8B HAT models trained on different datasets. On average, and for all but one individual benchmark comparison, Aleph-Alpha-GermanWeb outperforms FineWeb2. (Left) The HAT model was trained for 25,000 training steps on an English-language web-derived dataset, equating to $\sim$ 75 billion English words. Afterwards the model was further trained for 20,000 steps ($\sim$ 60 billion words) with German language data, either random FineWeb2 data augmented with high quality datasets like Wikipedia (FineWeb2:HQ Curated) or our generated synthetic data (Synthetic). Both datasets amount to 60 billion German words each. (Middle) We trained for 21,000 steps ($\sim$ 63 billion words) on either random FineWeb2 data or on our synthetic German data. (Right) The FineWeb2 training run was continued to 50,$0$ steps and is compared to a training of 50,000 steps on data from our Filtered CC pipeline. Both runs each equate to $\sim$ 150 billion words.
4 Discussion and Conclusions
@@
This paper presents a comprehensive data curation pipeline for German-language LLM pre-training, resulting in the creation of our high-quality German dataset, Aleph-Alpha-GermanWeb. By leveraging a combination of Common Crawl web data, FineWeb2, and synthetic data conditioned on organic web data, we demonstrate the effectiveness of our data curation pipeline and synthetic data generation in enhancing pre-training datasets. We trained two different model architectures on Aleph-Alpha-GermanWeb and FineWeb2 on a range of German-language benchmarks. Our results show that our dataset outperforms FineWeb2, even when combined with high-quality human-curated data sources, highlighting the potential of further iterating on existing data curation pipelines and synthetic data generation in advancing German-language LLM pre-training.
We believe there are many improvements and further research directions which will be important in this area. For example, we expect high-quality translation of existing, high-performing English datasets – such as explored in Wang, Lu, et al. [50] – to be a helpful augmentation method. In our manual inspection of machine-translated texts from English to German, however, we found a high number of errors. We therefore caution against training on large amounts of machine-translated data without also rigorously assessing its quality and naturalness.
Relatedly, a more thorough analysis of synthetic data and its apparent quality is needed. Recent work has demonstrated that recursively training models exclusively on synthetic data leads to a phenomenon termed ‘model collapse’ [43], wherein model performance gradually deteriorates over successive generations of training and synthesis. However, mixing synthetic data with organic data appears to avoid this issue [14]. Still, it remains unclear to what degree practitioners should opt for organic rather than synthetic data, or whether synthetic data is merely ‘gaming’ popular evaluation methods by providing data in a more relevant syntax than organic data.
@@
We make Aleph-Alpha-GermanWeb publicly available to the research community, with the hope that it will contribute to further advancements in this field. Our findings support the growing body of evidence on the importance of data curation and synthetic data generation in LLM pre-training, and we believe that our work will have a positive impact on the development of German-language LLMs.
Acknowledgments and Disclosure of Funding
We wish to thank: Johannes Messner, Carina Kauf, and Koen Oostermeijer for additional German language support; members of Aleph Alpha Research who assisted in the development and execution of model training and evaluation; members of the infrastructure support team; the FineWeb2 team for providing an excellent starting point; Common Crawl for freely providing such a valuable resource to the research community; and the NeMo Curator team for releasing their open-source curation tool.
Author contributions
Here we identify the main work-streams associated with this project and list the contributing authors in alphabetical order.
Data preparation: Michael Barlow, Bastian Harren, Volker Stampa, Stephan Wäldchen, Gregor     Ziegltrum
Common Crawl pipeline: Michael Barlow, Thomas F Burns, Bastian Harren, Letitia Parcalabescu,     Volker Stampa
Synthetic data generation: Michael Barlow, Thomas F Burns, Björn Deiseroth
Model-based filtering: Thomas F Burns, Bastian Harren, Letitia Parcalabescu, Stephan Wäldchen,     Gregor Ziegltrum
Quality-based ensembling: Michael Barlow, Letitia Parcalabescu
Experiments and evaluations: Björn Deiseroth
Manuscript writing and editing: Thomas F Burns, Letitia Parcalabescu, Stephan Wäldchen
Research and experimental design: Thomas F Burns, Björn Deiseroth
@@ Appendix A Prompt for LLM-as-a-judge in quality filtering
Educational filter – LLM-as-a-judge prompt SYSTEM: Deine Aufgabe ist es zu bewerten, wie gut die deutsche Sprachqualität eines deutschen Textauszugs ist. Gebe deine Bewertung in folgendem JSON-Format:
@@
     3: Es gibt sprachliche Schwächen, aber der Sinn kann gut erfasst werden.
@@ Appendix B Prompts for synthetic data generation
Summarization – data generation prompt Ihre Aufgabe ist es, den bereitgestellten Text gemäß diesen Anweisungen zu lesen und umzuformulieren: - Versuchen Sie, eine komprimierte, aber genaue und informative Version des Originaltextes zu erstellen, keine vereinfachte Zusammenfassung. - Erfassen und bewahren Sie die entscheidenden Informationen, Schlüsselkonzepte, wichtigen Werte und sachlichen Details im Originaltext auf und machen Sie ihn gleichzeitig lesbarer und zugänglicher. - Behalten Sie Fachbegriffe, Fachvokabular und komplexe Konzepte bei. - Bewahren Sie Beispiele, Erklärungen zu Denkschritten und unterstützende Beweise auf, um die Tiefe und den Kontext des Textes zu erhalten. - Fügen Sie nur Informationen hinzu, die im Originaltext vorhanden sind. Fügen Sie keine neuen oder unbegründeten Behauptungen hinzu. - Schreiben Sie im Klartext Beginne deine Antwort mit ’Umformulierte Version:’. Text: {document} Umformulierte Version:

### v3

@@ 2.3 Quality classification
Following prior work on quality classification for English-language web data Su et al. (2024a), we categorised documents into five quality buckets – high, medium-high, medium, medium-low, and low – by leveraging the strengths of several imperfect quality classifiers to produce a more accurate outcome. Specifically, we created several individual quality classifiers based on fastText Bojanowski et al. (2016) and Bidirectional Encoder Representations from Transformers (BERT) Devlin et al. (2019) models. In the following, we describe (1) our two grammar classifiers, (2) our two educational quality classifiers, (3) our two instruction classifiers and (4) the way in which we ensembled all classifiers and grouped the data into the five quality buckets according to the classifiers’ scores.
@@ 2.3.1 Grammar
Model Condition Points
Initialise all documents 0
BERT (edu. quality) Predicted score = 5 +3
fastText (edu. quality) Predicted “high quality” with >99% confidence +2
BERT (grammar) Predicted “high quality” +3
fastText (grammar) Predicted “high quality” with >99% confidence +2
BERT (instruction) High quality probability is in top 15% of the distribution +6
fastText (instruction) High quality probability is in top 15% of the distribution +4
@@ 2.3.2 Educational quality
Inspired by successful demonstrations for English with FineWeb Penedo et al. (2024) and Nemotron-CC Su et al. (2024a) (Apachee 2.0), we created educational quality classifiers to curate high-quality German web data by training a classifier on scores given by an LLM-as-a-judge. The intuition behind this is to identify web documents that are more likely to be educational, informative, and useful for model training, as opposed to spam, low-quality, or otherwise unhelpful content.
@@
For each document, we calculated a combined educational quality score by taking the minimum over the three criteria rated by the LLM-as-a-judge. This is because we observed that the judge would often miss poor quality in one criterium, but rarely in all three. We then used these scores as the training signal for fastText and BERT quality classification models. We trained a BERT model tasked to predict the scores given the first 512 tokens of the document’s text, and a binary fastText classifier.
@@ 2.3.3 Instruction
Building on Messmer et al. (2025), we trained binary classifiers using both fastText and BERT to distinguish between chat/instruction-style data (high quality) and other types of documents (low quality). To create a balanced training dataset, we created two subsets of 80,000 documents each. The low-quality subset consisted of randomly selected documents from the German FineWeb2 corpus, while the high-quality subset comprised randomly sampled instances from the Aya collection, Include Base-44, OpenAssistant2, and MMLU. To create samples from MMLU, we first concatenated a question and the respective answers. For OpenAssistant-2, we concatenated all samples that belong to the same conversation. We used 95% of the data for training and the remaining 5% for validation.
@@ 2.3.4 Ensembling
To split the data into different quality buckets, we use the rules in Table 3 to allocate overall quality points to each document. Using these scores, we categorised documents into five quality buckets based on the overall quality, as shown in Table 4.
Bucket Score Quality Distribution
@@
Medium-high 9119-11 15.0% 9.9%
Medium 585-8 36.3% 28.7%
Medium-low 343-4 15.5% 21.7%
Low <3<3 21.0% 31.6%
Table 4: FineWeb2 document quality classifications decided by overall quality score.
@@ 3 Experiments
We trained and evaluated a one billion parameter (1B) Llama-style transformer model and an eight billion parameter (8B) hierarchical autoregressive transformer (HAT) model on all three subsets of our dataset in isolation to assess their quality compared to FineWeb2.
@@ 3.1 Evaluations
The Multilingual Massive Multitask Language Understanding dataset Hendrycks et al. (2020), is a benchmark designed to evaluate the performance of LLMs across multiple languages and disciplines. It extends the original MMLU benchmark by translating its test set into 14 languages—including German—using professional human translators to ensure accuracy. The dataset encompasses 57 subjects ranging from elementary-level topics to advanced professional fields such as law, physics, history, and computer science. The German portion contains 14,000 samples.
The following three popular LLM benchmarks were translated from English to German Plüster .
@@
The AI2 Reasoning Challenge (ARC) Clark et al. (2018) is a benchmark dataset developed by the Allen Institute for AI to evaluate the reasoning capabilities of artificial intelligence models. It comprises 7,787 multiple-choice science questions sourced from standardized tests designed for students in grades three through nine. The dataset is divided into two subsets: ARC-Easy and ARC-Challenge; here we evaluate the translation of ARC-Easy.
@@
The HellaSwag benchmark dataset Zellers et al. (2019) is designed to evaluate the commonsense reasoning abilities of AI models, particularly in the context of sentence completion tasks. It comprises approximately 70,000 multiple-choice questions from diverse sources, including instructional videos and articles from platforms like WikiHow and ActivityNet. Each question presents a context followed by four possible sentence completions, one of which is correct.
@@
The TruthfulQA dataset Lin et al. (2021) is a benchmark designed to evaluate the truthfulness of language models when generating answers to questions. It comprises 817 questions across 38 categories, including health, law, finance, and politics. The questions are crafted to challenge models with scenarios where humans might hold incorrect beliefs or misconceptions, aiming to assess whether models can avoid generating false answers learned from imitating human texts.
Figure 1: Average accuracies (MMMLU, ARC-Easy, HellaSwag; single- & five-shot) of 1B Llama-style models trained on $\sim$ 84 billion tokens from different datasets. All subsets of GermanWeb – marked with an asterisk (*) – outperform FineWeb2. For comparison, we also include EPFML High. Here, 1,000 training steps equates to approximately 2.1 billion tokens. Individual benchmark results are provided in Appendix D.1.
@@
We trained and evaluated our filtered Common Crawl (Filtered CC), synthetic, and our FineWeb2 High bucket, and compared these to random sampling German FineWeb2 and the top 10% of German FineWeb2 documents classified by Messmer et al. (2025) (EPFML High). The avearge accuracies across all benchmarks are shown in Figure 1 and the per-benchmark breakdowns are shown in Appendix D.1. We chose to exclude TruthfulQA from the 1B parameter model evaluations, as we found that it did not give a clear signal and suspect that this size of the model may be incapable of the task (cf. §3.3 for results of an 8B model).
We wish to emphasise that a large part of the synthetic subset’s average performance gain stems from its far better performance on the MMLU task. Its performance on ARC is still better than FineWeb2, but not clearly the best performer. And for the Hellaswag benchmark, while it again outperforms FineWeb2, high-quality organic data appear even better. This suggests that mixing differently-derived datasets may be appropriate to balance strengths and weaknesses of individual subsets.
@@
To evaluate our dataset on a different architecture and a larger scale, we pre-trained 8B hierarchical autoregressive transformer (HAT) models Neitemeier et al. (2025a) (see Appendix C.2 for training details). With this we leverage recent advances demonstrating that jointly training the tokeniser and model can be a promising alternative Pagnoni et al. (2024), aligning with the philosophy of end-to-end learning  Sutton (2019).
An additional benefit of using an end-to-end, tokeniser-free architecture like HAT is that it debiases comparisons between datasets—comparisons which are a central focus of this paper. Tokenisers have a significant impact on the efficacy of model training Ali et al. (2024); Yang et al. (2024); Wang et al. (2024). Typically they are trained on large text corpora, which makes models using those tokenisers more effective on datasets that resemble the tokeniser’s training distribution. As a result, using a tokeniser-based model can introduce bias when comparing the effectiveness of different pre-training datasets, favouring those models that align more closely with the tokeniser’s original training data Mayilvahanan et al. (2025). To avoid this bias, we use a HAT model, which does not rely on any pre-trained tokeniser and instead performs fixed, dataset-agnostic, byte-level string splitting.
@@
When pre-training LLMs, web data is often mixed from different languages and with data from alternative, curated sources Grattafiori et al. (2024), e.g., English and German data might be mixed with additional non-web data from books and encyclopedias. To test the performance of our single highest-performing subset, we compared our synthetic data to mix of 50% FineWeb2 and 50% high-quality curated datasets [^5] In particular: the German National Library, European Union Parliament transcripts, European Patent Office applications, German Text Archive, Project Gutenberg books, JRC-Acquis, NewsAPI, Open Legal Data, Open Web Math, peS2o, PhilPapers, Wikibooks, Wikinews, Wikpedia, Wikisource, and Wikivoyage.. Figure 2 (left) shows our synthetic data outperformed this mixture on our tested benchmarks.
@@
Figure 2: Accuracies (0-shot) of 8B HAT models trained on different datasets. On average, and for all but one individual benchmark comparison, GermanWeb subsets – marked with an asterisk (*) – outperform FineWeb2. (Top) We trained for 25,000 training steps on an English-language web-derived dataset, equating to $\sim$ 75 billion English words. Afterwards the model was further trained for 20,000 steps ($\sim$ 60 billion words) with German language data, either random FineWeb2 data augmented with high quality datasets like Wikipedia (FineWeb2:HQ Curated), and listed further in Appendix C.3, or our generated synthetic data (Synthetic). Both datasets amount to 60 billion German words each. (Middle) We trained for 21,000 steps ($\sim$ 63 billion words) on either random FineWeb2 data or on our synthetic German data. (Bottom) The FineWeb2 training run was continued to 50,$00$ steps and is compared to a training of 50,000 steps on data from our Filtered CC pipeline. Both runs each equate to $\sim$ 150 billion words.
4 Conclusion
@@
This paper presents a comprehensive data curation pipeline for German-language LLM pre-training, resulting in the creation of our high-quality German dataset, GermanWeb composed of three subsets drawn from Common Crawl, FineWeb2, and synthetic data conditioned on organic web documents. We evaluate each subset in isolation and benchmark them against FineWeb2 training demonstrating the effectiveness of our data curation pipeline and synthetic data generation in enhancing pre-training datasets. We trained two different model architectures on GermanWeb and FineWeb2 on a range of German-language benchmarks. Our results show that all three subsets of our dataset individually outperform FineWeb2, even when the latter is combined with high-quality human-curated data sources. This highlights the potential of future work (i) iterating on existing data curation pipelines, (ii) improving synthetic data generation, and (iii) mixing organic and synthetic data for advancing German-language LLM pre-training.
@@
We make GermanWeb publicly available to the research community to contribute to further advancements in this field. Our findings support the importance of data curation and synthetic data generation in LLM pre-training, and we believe that our work will have a positive impact on the development of German-language LLMs.
5 Limitations
We believe there are many improvements and further research directions which will be important in this area. For example, we expect high-quality translation of existing, high-performing English datasets – such as explored in Wang et al. (2025) – to be a helpful augmentation method. In our manual inspection of machine-translated texts from English to German, however, we found a high number of errors. We therefore caution against training on large amounts of machine-translated data without also rigorously assessing its quality and naturalness.
Relatedly, a more thorough analysis of synthetic data and its apparent quality is needed. Recent work has demonstrated that recursively training models exclusively on synthetic data leads to a phenomenon termed ‘model collapse’ Shumailov et al. (2024), wherein model performance gradually deteriorates over successive generations of training and synthesis. However, mixing synthetic data with organic data appears to avoid this issue Gerstgrasser et al. (2024). Still, it remains unclear to what degree practitioners should opt for organic rather than synthetic data, or whether synthetic data is merely ‘gaming’ popular evaluation methods by providing data in a more relevant syntax than organic data.
@@ Appendix A Prompt for LLM-as-a-judge in quality filtering
Educational filter – LLM-as-a-judge prompt (1/2) SYSTEM: Deine Aufgabe ist es zu bewerten, wie gut die deutsche Sprachqualität eines deutschen Textauszugs ist. Gebe deine Bewertung in folgendem JSON-Format:
@@
Educational filter – LLM-as-a-judge prompt (2/2) continued 3: Es gibt sprachliche Schwächen, aber der Sinn kann gut erfasst werden.
@@ Appendix B Prompts for synthetic data generation
Summarisation – data generation prompt Ihre Aufgabe ist es, den bereitgestellten Text gemäß diesen Anweisungen zu lesen und umzuformulieren: - Versuchen Sie, eine komprimierte, aber genaue und informative Version des Originaltextes zu erstellen, keine vereinfachte Zusammenfassung. - Erfassen und bewahren Sie die entscheidenden Informationen, Schlüsselkonzepte, wichtigen Werte und sachlichen Details im Originaltext auf und machen Sie ihn gleichzeitig lesbarer und zugänglicher. - Behalten Sie Fachbegriffe, Fachvokabular und komplexe Konzepte bei. - Bewahren Sie Beispiele, Erklärungen zu Denkschritten und unterstützende Beweise auf, um die Tiefe und den Kontext des Textes zu erhalten. - Fügen Sie nur Informationen hinzu, die im Originaltext vorhanden sind. Fügen Sie keine neuen oder unbegründeten Behauptungen hinzu. - Schreiben Sie im Klartext Beginne deine Antwort mit ’Umformulierte Version:’. Text: {document} Umformulierte Version: